Define all of your unknowns and use at least 2 distinct variables for the problem.
The sum of 3 integers is 120. The second integer is double the first. The third is triple the first. Find the integers.

Answer by Boreal(15235)

integers are x, 2x, third is y
x+2x+y=120
3x+y=120
y=120-3x
y=3x
therefore, 3x=120-3x
6x=120
x=20
2x=40
120-3x=60
numbers are 20, 40, 60